OpenStudy (anonymous):

find all real solutions of the equation 19. (3x+5)^4 - (3x+5)^3=0

OpenStudy (anonymous):

\[=\left( 3x+5 \right)^3\left( 3x+5-1 \right)=0\] \[=\left( 3x+5 \right)^3\left( 3x+4 \right)=0\] ?

OpenStudy (anonymous):

find x